# Robust Learning in Bayesian Parallel Branching Graph Neural Networks: The Narrow Width Limit

### v2

@@ 5 BPB-GNN on Cora
An interesting aspect of the BPB-GNN network is that the branch readout norms converge at the narrow width for different hyperparameters $\sigma_{w}$ and $L$, reflecting the natural branch importance for the task.
Figure 6: Cora experiment: statistical average of squared readout norms $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ for each branch $l$ as a function of the network width $N$, and regularization strength $\sigma_{w}$. The BPB-GNN has $L=6$ branches.
As shown in Figure 6, the BPB-GNN with branches $L=6$ robustly learns the readout norms at narrow width independently of $\sigma_{w}$’ s, consistent with the student-teacher results. This suggests that we can recast the data as generated from a ground-truth teacher network even for real-world datasets. The last branch of the BPB-GNN network has a larger contribution, reflecting the presence of higher-order convolutions in the Cora dataset. From a kernel perspective, increasing branches better distinguish the nodes, as shown in Figure 7. This could explain the selective turn-off of intermediate branches and the increased contribution of the last branch.
@@
Furthermore, the first two branches are learned most robustly at narrow width, as shown in Figure 8, where the branch norms converge for the first two branches even for BPB-GNNs with different $L$. This suggests that the branch importance, as reflected by the norms learned at narrow width, indicates the contribution of the bare data and the first convolution layer.
@@ 6 Discussion
The findings presented in this paper reveal that BPB-GNNs exhibit unique characteristics in the narrow width limit. Unlike the infinite-width limit, where neural networks behave as Gaussian Process (GP) with task-independent kernels, narrow-width BPB-GNNs undergo significant kernel renormalization. This renormalization leads to breaking of the symmetry between the branches, resulting in more robust and differentiated learning. Our experiments demonstrate that narrow-width BPB-GNNs can retain and, in some cases, improve generalization performance compared to their wider counterparts, particularly in bias-limited scenarios where the regularization effects dominate. Additionally, the observed independence of readout norms from architectural hyperparameters suggests that narrow-width BPB-GNNs can capture the intrinsic properties of the data more effectively. These insights suggest potential new strategies for optimizing neural network architectures in practical applications, challenging the traditional emphasis on increasing network width for better performance. Although our work is focused on GNN architectures, our findings can be transferred to the transformer family of architectures, as the transformer with frozen attention is analogous to a fully connected GNN (Vaswani et al. (2017); Veličković et al. (2017)), and a Bayesian theory with kernel renormalization can be developed similarly as in Tiberi et al. (2024).
@@
We demonstrate that the bias term robustly decrease at narrow width; however, the variance term sometimes dominates and it is not clear BPB-GNN generalization is provably smaller at narrow width. Furthermore, our results rely on the Bayesian network setting, which might not directly transfer to understanding generalization of GNN trained with SGD, for which the loss landscape may be different.
@@ 7 Conclusion
In conclusion, this paper introduces and investigates the concept of narrow width limits in Bayesian Parallel Branching Graph Neural Networks. Contrary to the common belief that wider networks inherently generalize better, our results indicate that BPB-GNNs with significantly narrower widths can achieve better or competitive performance. This is attributed to effective symmetry breaking and kernel renormalization in the narrow-width limit, which leads to robust learning. Our theoretical analysis, supported by empirical evidence, establishes a new understanding of how network architecture influences learning outcomes. This work provides a novel perspective on neural network design and suggests further research into optimizing network structures for specific learning tasks, especially in scenarios where model simplicity and computational efficiency are crucial.
@@ acknowledgement
Appendix A Details on Theory of BPB-GNN
@@ A.1 Summary of Notations
$N$ BPB-GNN hidden layer width
@@ A.2 Kernel renormalization
First, we linearize the energy in terms of $a_{l}$’s by introducing the auxiliary variables $t^{\mu},\mu=1,\dots,P$.
@@
Now we perform the integration on $W_{l}$’s, and get a Fourier representation of $Z$ with $h_{l}$,$u_{l}$ as auxiliary variables after inserting $t$:
$Z$ $=\int\prod_{l=0}^{L}dh_{l}du_{l}dt\exp\Big{(}it^{T}Y-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}-\frac{1}{2}t^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}t\Big{)}$
$=\int\prod_{l=0}^{L}dh_{l}du_{l}\exp\Big{(}-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}\frac{1}{2}Y^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}^{-1}Y\Big{)}$ (29)
@@ A.4 Statistics of branch readout norms
Appendix B Experimental details
B.1 Student-teacher CSBM
@@
In the experiment, we use $N_{0}=950$,$d=20$,$\lambda=4$ and $\mu=4$. The teacher network parameters are variance $\sigma_{t}^{2}=1$, width $N_{t}=1024$, branch norms variance $\beta_{0}^{2}=0.4$, $\beta_{1}^{2}=2$ for individual element of the readout vector $a_{l}$. Temperature $T=0.0005\sigma_{w}^{2}$ for each $\sigma_{w}$ value.
B.2 Cora
For the Cora dataset, we use a random split of the data into 21% as training set and 79% as test set. We group the classes $(1,2,4)$ into one group and the rest for the other group for binary node regression, with labels as ±$\pm 1$’s. The Bayesian theory and HMC sampling follows the same design as in the student-teacher setup. We use temperature $T=0.01$ for both theory and sampling as the sampling becomes more difficult for smaller temperature. This explains the discrepancy of the GP limit bias for different $\sigma_{w}$ values.
B.3 Hamiltonian Monte Carlo
The sampling experiments in the paper are all done with Hamiltonian Monte Carlo (Betancourt (2017); Neal (2012)) simulations, a popular method for sampling a probability distribution. HMC has faster convergence to the posterior distribution compared to Langevin dynamics. We used Numpyrho to set up chains and run the simulations on the GPU cluster. Due to memory constraint, we only sampled up to $N=1024$ hidden layer width for the student-teacher CSBM experiment and $N=64$ for the Cora experiment. Since we mainly aim to demonstrate the narrow width effect in this paper, this suffices the purpose.

### v3

@@ 5 BPB-GCN on Cora
An interesting aspect of the BPB-GCN network is that the branch readout norms converge at the narrow width for different hyperparameters $\sigma_{w}$ and $L$, reflecting the natural branch importance for the task.
Figure 6: Cora experiment: statistical average of squared readout norms $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ for each branch $l$ as a function of the network width $N$, and regularization strength $\sigma_{w}$. The BPB-GCN has $L=6$ branches.
As shown in Figure 6, the BPB-GCN with branches $L=6$ robustly learns the readout norms at narrow width independently of $\sigma_{w}$’ s, consistent with the student-teacher results. This suggests that we can recast the data as generated from a ground-truth teacher network even for real-world datasets. The last branch of the BPB-GCN network has a larger contribution, reflecting the presence of higher-order convolutions in the Cora dataset. From a kernel perspective, increasing branches better distinguish the nodes, as shown in Figure 7. This could explain the selective turn-off of intermediate branches and the increased contribution of the last branch. Our results also indicate that there is no oversmoothing problem at narrow width: as $L$ increases, individual branches can still learn robustly.
@@
Furthermore, the first two branches are learned most robustly at narrow width, as shown in Figure 8, where the branch norms converge for the first two branches even for BPB-GCNs with different $L$. This suggests that the branch importance, as reflected by the norms learned at narrow width, indicates the contribution of the bare data and the first convolution layer.
@@ 6 Discussion
The findings presented in this paper reveal that BPB-GCNs exhibit unique characteristics in the narrow width limit. Unlike the infinite-width limit, where neural networks behave as Gaussian Process (GP) with task-independent kernels, narrow-width BPB-GCNs undergo significant kernel renormalization. This renormalization leads to breaking of the symmetry between the branches, resulting in more robust and differentiated learning. Our experiments demonstrate that narrow-width BPB-GCNs can retain and, in some cases, improve generalization performance compared to their wider counterparts, particularly in bias-limited scenarios where the regularization effects dominate. Therefore, there is a trade-off between the expressive power of GCN Xu et al. (2018); Loukas (2019) and stability Stogin et al. (2020). Additionally, the observed independence of readout norms from architectural hyperparameters suggests that narrow-width BPB-GCNs can capture the intrinsic properties of the data more effectively. These insights suggest potential new strategies for optimizing neural network architectures in practical applications, challenging the traditional emphasis on increasing network width for better performance. Although our work is focused on GCN architectures, our findings are more general for a family of branching networks as shown in Appendix B, and in particular the Bayesian theory with kernel renormalization is shown to be successful for CNN Aiudi et al. (2023) as well as transformers Tiberi et al. (2024). Intriguingly, Vyas et al. (2024) finds that an ensemble of ResNets averaged over random initializations accidentally shows this narrow width effect, for which our findings provide an explanation.
@@
We demonstrate that the bias robustly decreases at narrow width for the branching networks due to kernel renormalization, which defies the common belief that wider networks are better (Allen-Zhu et al. (2019); Lee et al. (2017)). However, we do find that for real-world scenarios such as BPB-GCN on Cora, and residual-MLP on Cifar10, the generalization error is usually variance-limited. In principle, training an ensemble of branching networks averaged over random initialization helps to reduce the variance significantly such that the narrow-width effect is more pronounced as in Vyas et al. (2024). Furthermore, since our Bayesian setup corresponds to an ensemble of fully trained over-parametrized networks in an offline fashion, the learning is still in the so-called lazy regime with minimal feature learning (Atanasov et al. (2022); Karkada (2024)). We hypothesize that narrower width serves as a regularizer in this regime which might break down as the network is trained with online learning and approaches the rich regime (Yang et al. (2022)) with more feature learning. In that case, it might be possible that even the bias does not decrease at the narrow width.
@@ 7 Conclusion
In conclusion, this paper introduces and investigates the concept of narrow width limits in Bayesian branching networks. Our results indicate that branching networks with significantly narrower widths can achieve better or competitive performance, contrary to common beliefs. This is attributed to effective symmetry breaking and kernel renormalization in the narrow-width limit, which leads to robust learning. Our theoretical analysis, supported by empirical evidence, establishes a new understanding of how network architecture influences learning outcomes. This work provides a novel perspective on the infinite width limit of neural networks and suggests further research directions in understanding narrow width as a regularization effect.
@@ acknowledgement
Appendix A Details on Theory of BPB-GCN
@@ A.1 Summary of Notations
$N$ BPB-GCN hidden layer width
@@
$\langle x\rangle$ Statistical average of any quantity $x$ over the posterior distribution
@@ A.2 Kernel renormalization
First, we linearize the energy in terms of $a_{l}$’s and transform the integral by introducing the auxiliary variables $t^{\mu},\mu=1,\dots,P$.
@@
Now we perform the integration on $W_{l}$’s and introduce the auxiliary variables $t^{\mu}$’s again using the Gaussian trick and get
[…]
where $G(t)$ is in terms of the kernel averaged over the Gaussian measure
$G(t)=\log\left\langle\exp\left(-\frac{1}{2N}t^{T}K(W)t\right)\right\rangle_{W}$ (28)
Writing out the integral explicitly, we have
$G(t)=\log\int\prod_{l=0}^{L-1}DW_{l}\exp\left(-\frac{1}{2N}\sum_{\mu,\nu}t^{\mu}t^{\nu}\sum_{j,l}\frac{\sigma_{w}^{2}}{N_{0}L}(\sum_{\mu^{\prime},i}A_{\mu,\mu^{\prime}}^{l}W_{ij}^{l}x_{i}^{\mu^{\prime}})(\sum_{\nu^{\prime},i^{\prime}}A_{\nu,\nu^{\prime}}^{l}W_{i^{\prime}j}^{l}x_{i^{\prime}}^{\nu^{\prime}}))\right)$ (29)
where $DW_{l}$ is the Gaussian measure
$DW_{l}=\prod_{i,j}e^{-W_{l,i,j}^{2}/2\sigma_{w}^{2}}dW_{l,i,j}$ (30)
By recognizing that this is just a product of the same $N$ integrals on $N_{0}$ dimensional weight vector $\vec{W}^{j}_{l}$ for a fixed $j$ neuron, we get
$G(t)=N\log\int\prod_{l=0}^{L-1}d\vec{W}^{j}_{l}\exp\left(\sum_{l,i,i^{\prime}}(-\frac{1}{2N}\sum_{\mu,\nu}t^{\mu}t^{\nu}\frac{\sigma_{w}^{2}}{N_{0}L}\xi_{l,i}^{\mu}\xi_{l,i^{\prime}}^{\nu}+\frac{1}{2\sigma_{w}^{2}}\delta_{i,i^{\prime}})W^{j}_{i,l}W^{j}_{i^{\prime},l}\right)$ (31)
where $\xi_{l,i}^{\mu}=\sum_{\mu^{\prime}}A^{l}_{\mu,\mu^{\prime}}x_{i}^{\mu^{\prime}}$ is the $l$th branch convolution on node $\mu$. Since the network is linear, we are able to perform the Gaussian integration exactly and get
$G(t)=-\frac{N}{2}\sum_{l}\log(\det(I_{i,i^{\prime}}+\frac{\sigma_{w}^{2}}{N_{0}L}t^{\mu}\xi^{\mu}_{i,l}\xi^{\nu}_{i^{\prime},l}t^{\nu}))$ (32)
where the determinant is over the $i,i^{\prime}$ matrix, and thus performing the determinant we arrive at
$G(t)=-\frac{N}{2}\sum_{l}\log(1+t^{T}\frac{K_{l}}{L}t)),$ (33)
where $K^{l}_{\mu,\nu}=\frac{\sigma_{w}^{2}}{N}\sum_{i}t^{\mu}\xi^{\mu}_{i,l}\xi^{\nu}_{i,l}$ is the Gaussian process kernel for branch $l$.
Next, we insert $L$ delta functions $\int\prod_{l}dh_{l}\delta(h_{l}-t^{T}K_{l}t)$ to enforce the identity $G(t)=\sum_{l}-\frac{N}{2}\log(1+h_{l})$ and use the Fourier representation of it with auxiliary variables $u_{l}$ to get:
$Z$ $=\int\prod_{l=0}^{L-1}dh_{l}du_{l}dt\exp\Big{(}it^{T}Y-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}-\frac{1}{2}t^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}t\Big{)}$
$=\int\prod_{l=0}^{L-1}dh_{l}du_{l}\exp\Big{(}-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma_{w}^{2}}u_{l}h_{l}\frac{1}{2}Y^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}^{-1}Y\Big{)}$ (34)
@@ A.4 Statistics of branch readout norms
Appendix B The narrow width limit for general architectures
In this section, we demonstrate that our results on the narrow-width limit can be generalized to any network with branching structures, supported by both theoretical and empirical evidence.
Consider a general branching network, with $L$ independent branches each with hidden layer weight $W_{l}$ and readout weight $a_{l}$, ie.
$f(x;\Theta)=\sum_{l=0}^{L-1}\frac{1}{\sqrt{L}}f_{l}(x;\Theta_{l})=\sum_{l}\frac{1}{\sqrt{NL}}a_{l}\phi_{l}(x;W_{l}),$ (60)
where $\phi_{l}(x;W_{l})$ represents any feed-forward function on the dataset $X$, ie. in matrix form the branch $l$ readout is $F_{l}=\phi_{l}(\frac{1}{\sqrt{N_{0}}}XW_{l})$. As an example, $\phi_{l}(XW_{l})=\frac{1}{\sqrt{N_{0}}}A^{l}XW_{l}$ represents $l$ convolutions on the node features in the BPB-GCN architecture. $\phi_{l}(XW_{l})=\frac{1}{\sqrt{N_{0}}}A_{l}XW_{l}$ can also represent an attention head $l$ for attention-networks with frozen attention $A_{l}$, as well as a kernel convolution for patch $l$ in the CNN architecture. In the residual-MLP example below, $\phi_{l}$ represents either the linear or the ReLu branch.
With the general branching network, we can similarly consider the posterior distribution in the Bayesian setup as Eq. 4 with squared error as the likelihood term and $L_{2}$ regularization as the prior term. Provided the network is linear, the partition function Eq.5 can be integrated in a similar fashion and arrive at the saddle point equation
with $K=\sum_{l}u_{l}K_{l}/L$ and $K_{l}$ corresponding to the NNGP kernel (Lee et al. (2017)). In fact, even with non-linear activations, we conjecture that the kernel renormalization still holds with the order parameters $u_{l}$ that satisfies the saddle point equation, and the posterior weight statistics calculated with the renormalized kernel agrees with the experiments as empirically shown in the residual-MLP section.
B.1 The generalized equipartition theorem
With the general branching network, we show the narrow width limit result for the student-teacher setup. Specifically we prove the following
The generalized equipartition theorem
: For the general branching network Eq. 60, where the teacher network width is $N_{t}$ with $W_{l,ij}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $a^{*}_{i,l}\sim\mathcal{N}(0,\beta_{l}^{2})$ and the student network with layer width $N$ in the Bayesian regression setup of Eq. 4, at $T=0$ temperature, in the limit $N_{t}\to\infty$ and $\alpha=P/N\to\infty$, there exists a solution to the saddle point equations 12
$u_{l}=\sigma_{t}^{2}\beta_{l}^{2}/\sigma_{w}^{2}.$ (62)
Proof: The proof is exactly the same as before, except for changing $K_{l}$ as a general NNGP kernel that corresponds to the feature map $\phi_{l}$ for branch $l$.
B.2 Residual-mlp
In the following, we show a simple 2-layer residual network as an example for the general branching networks, with a linear branch $\phi_{0}(x)=\frac{1}{\sqrt{N_{0}}}W_{0}x$ and a ReLU branch $\phi_{1}(x)=\frac{1}{\sqrt{N_{0}}}\mathrm{ReLu}(W_{1}x)$. Therefore, the overall kernel is $K=\frac{1}{2}(u_{0}K_{0}+u_{1}K_{1})$, with the NNGP kernels $K_{l}$ given by
$K_{0}=K_{linear}=\frac{\sigma^{2}}{N_{0}}X_{0}X_{0}^{T}$ (63)
and the relu kernel (Cho & Saul (2009))
$K_{1}(x,x^{\prime})=K_{relu}(x,x^{\prime})=\frac{1}{2\pi}\sqrt{K_{0}(x,x)K_{0}(x^{\prime},x^{\prime})}J(\theta),$ (64)
where
$J(\theta)=\sin(\theta)+(\pi-\theta)\cos(\theta)$ (65)
and
$\theta=\cos^{-1}(\frac{K_{0}(x,x^{\prime})}{\sqrt{K_{0}(x,x)K_{0}(x^{\prime},x^{\prime})}})$ (66)
For the student-teacher task (details in Appendix C.3 ), we observe the same narrow-width limit phenomenon as in the BPB-GCN case, where each student branch robustly learns the teacher’s corresponding branch but approaches the same GP limit as the network width gets larger.
Figure 9: Statistical average of student readout norms squared as a function of network width from theory and HMC sampling, for the residual-MLP student-teacher task, with fixed $\sigma_{w}=1$.
As shown in Figure 9 and 10, there is a symmetry breaking of branches from the GP-limit to the narrow-width limit, with the ReLu branch and the linear branch learning their corresponding teacher branch norms as $N$ decreases.
Refer to caption Refer to caption
(a) (b)
Figure 10: $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ as a function of network width $N$ for a range of $\sigma_{w}$ values. (a) Linear branch; (b) Relu branch. The student branch norms with different regularization strengths all converge to the same teacher readout norm values at narrow width.
Furthermore, narrower network performs better in terms of the bias and larger $\sigma_{w}$ corresponds to a shorter transition window from the GP-limit to the narrow-width limit, similar to the BPB-GCN case. As shown in Figure 11 and 12, the trade-off between the bias and variance contribution is more pronounced for small $\sigma_{w}$’s, usually with an optimal width for the generalization error. These results support our generalized equipartition theorem and demonstrates that the narrow width limit is a general phenomenon for branching networks.
Refer to caption Refer to caption
(a) (b)
Figure 11: Student residual-MLP network generalization performance as a function of network width $N$ and regularization strength $\sigma_{w}$. (a) Bias; (b) Variance.
Figure 12: Overall generalization as a sum of bias and variance in Figure 11.
B.3 Residual-mlp on Cifar10
We also perform theory and experiments on the Cifar10 (Krizhevsky et al. (2009)) dataset for a binary classification task (translated to Bayesian regression), with details elaborated in Appendix C.4. Similarly, there exists symmetry breaking of branches and the branch norms converge at the narrow width for different $\sigma_{w}$’s, as shown in Figure 13,14. Interestingly, the ReLu branch completely dominates at the narrow width, which means the ReLu branch is much more important for the residual-MLP when the training sample size is much bigger than the network width.
Figure 13: Statistical average of student readout norms squared as a function of network width from theory and HMC sampling, for the Cifar10 dataset with residual-MLP, with fixed $\sigma_{w}=1$.
Refer to caption Refer to caption
(a) (b)
Figure 14: Cifar10 $\langle\|a_{l}\|^{2}\rangle\sigma_{w}^{2}/N$ as a function of network width $N$ for a range of $\sigma_{w}$ values. (a) Linear branch; (b) Relu branch. The student branch norms with different regularization strengths all converge to the same teacher readout norm values at narrow width.
Furthermore, we also show in Figure 15,16(a) that the dominance of ReLu branch at narrow width drives the bias to decrease, albeit overall genralization is dominated by the variance. Using the average of the label predictors over different experimental samples, we show that the test accuracy indeed decreases with larger network width $N$ in Figure 16(b).
Refer to caption Refer to caption
(a) (b)
Figure 15: Residual-MLP generalization performance on Cifar10 as a function of network width $N$ and regularization strength $\sigma_{w}$. (a) Bias; (b) Variance.
Refer to caption Refer to caption
(a) (b)
Figure 16: (a) Overall generalization error for residual-MLP on Cifar10. (b) Test accuracy on Cifar10, calculated using the class label by averaging the predictions from HMC samples.
Appendix C Experimental details
C.1 Student-teacher CSBM
@@
In the experiment, we use $N_{0}=950$,$d=20$,$\lambda=4$ and $\mu=4$. The teacher network parameters are variance $\sigma_{t}^{2}=1$, width $N_{t}=1024$, branch norms variance $\beta_{0}^{2}=0.4$, $\beta_{1}^{2}=2$ for individual element of the readout vector $a_{l}$. We used total number of nodes $n=2600$ with train ratio 0.65 for semi-supervised node regression. Temperature $T=0.0005\sigma_{w}^{2}$ for each $\sigma_{w}$ value in $\sigma_{w}=\{0.5,0.8,1.0,1.2\}$.
C.2 Cora
For the Cora dataset, we use a random split of the data (total nodes $n$ =2708) into 21% as training set and 79% as test set. We group the classes $(1,2,4)$ into one group and the rest for the other group for binary node regression, with labels as ±$\pm 1$’s. The Bayesian theory and HMC sampling follows the same design as in the student-teacher setup. We use temperature $T=0.01$ for both theory and sampling as the sampling becomes more difficult for smaller temperature. This explains the discrepancy of the GP limit bias for different $\sigma_{w}$ values in $\sigma_{w}=\{1.0,1.2,1.4\}$
C.3 student-teacher residual-mlp
For the student-teacher task of the residual MLP, we use the data matrix $X$ generated from i.i.d unit Gaussian distribution, with input dimension $N_{0}=1024$, training sample size $P=1280$ and test sample size $P_{test}=320$. The teacher network weights are also generated with i.i.d Gaussian, with the hidden layer width $N_{t}=1024$, hidden layer weight variance $\sigma_{w}^{2}=1$, and readout branch variance $\beta_{l}^{2}=(2.4,0.4)$ for the two branches. The temperature is set as $T=0.001\sigma_{w}^{2}$ for a range of $\sigma_{w}$ values in $\sigma_{w}=\{0.6,0.8,1.0,1.2,1.5\}$.
C.4 Cifar10
For the Cifar10 dataset, we filter the category of airplane and that of cat for the regression task. We use a random split of training sample size $P=1000$ and test sample size $P_{test}=1000$. The input data is only from the red channel with input dimension $N_{0}=1024$. The temperature is set at $T=0.005\sigma_{w}^{2}$ for a range of $\sigma_{w}$ values in $\sigma_{w}=\{1.0,1.2,1.5\}$. Within the compute budget, we observe significant sample correlations, which has a large impact on measuring the bias term. This is why we observe relatively large deviation of the bias results from theory in Figure 15(a).
C.5 Hamiltonian Monte Carlo
The sampling experiments in the paper are all done with Hamiltonian Monte Carlo (Betancourt (2017); Neal (2012); Chandra et al. (2021)) simulations, a popular method for sampling a probability distribution. HMC has faster convergence to the posterior distribution compared to Langevin dynamics. We used Numpyrho to set up chains and run the simulations on the GPU cluster. All experiments are within 20 GPU hours budget on Nvidia A100-40G. Due to memory constraint, we only sampled up to $N=1024$ hidden layer width for the student-teacher CSBM experiment and $N=64$ for the Cora experiment. Since we mainly aim to demonstrate the narrow width effect in this paper, this suffices the purpose.